# ModeTv2: GPU-accelerated Motion Decomposition Transformer for Pairwise Optimization in Medical Image Registration

### v1

@@ 3.3.1 GPU-accelerated ModeT
Let $F,M\in\mathbb{R}^{c\times h\times w\times l}$ be the fixed and moving feature maps from a specific layer of the hierarchical encoder, where $h,w,l$ are feature map dimensions, and $c$ is the channel count. These feature maps undergo linear projection ($proj$) and LayerNorm ($LN$) operations (Ba et al., 2016) to generate $Q$ (query) and $K$ (key):
@@
where the shared-weight projection operation originates from $N(0,1e^{-5})$, with biases initialized to 0. $Q$ and $K$ are then divided based on channels, and $S$ represents the number of attention heads.
The neighborhood attention map is then calculated. Let $c(p)$ be the voxel $p$’s neighborhood. For a neighborhood of size $n\times n\times n$, $||c(p)||=n^{3}$. The multi-head neighborhood attention map is obtained through:
@@
where $B\in\mathbb{R}^{S\times n\times n\times n}$ is a learnable relative positional bias, initialized to zeros. Zero padding is applied to the moving feature map to handle boundary voxels. Equation (3) shows how the neighborhood attention is computed for the $s$-th head at position $p$, breaking down semantic information from low resolution for individual similarity computation, laying the foundation for modeling diverse motion modes.
@@
where $\varphi^{(s)}\in\mathbb{R}^{h\times w\times l\times 3}$, $V\in\mathbb{R}^{n\times n\times n}$, and $V$ (value) represents the relative position coordinates for the neighborhood centroid. Importantly, $V$ remains unlearned, allowing the multi-head attention relationship to naturally transform into a multi-coordinate association.
@@ 3.3.2 Registration Head (RegHead)
We reevaluate the challenges associated with the diverse motion modes obtained through ModeT. In (Wang et al., 2023), a competitive weighting module (CWM) is designed to select a predominant motion pattern from various motion modes at a single point. However, weighting alone is insufficient to adequately address issues such as local field crossings that arise from multiple subfields (see Fig. 3(a)), since ModeT does not account for neighborhood relationships during calculation, or more precisely, only considers point to-point relationships. Simultaneously, to achieve a highly generalized network, we aim to minimize the number of trainable parameters on the deformation field estimator. We prefer the network to focus more on relearning the encoder’s feature extraction during PO.
@@
Under the condition of employing the diffeomorphic layer, we define the output of the convolution layer in RegHead as the stationary velocity field $v$ (Dalca et al., 2019). With regard to time $t$, the deformation field $\phi$ is defined as the ordinary differential equation (ODE):
@@
where $\phi^{(t)}$ is an identity transformation when $t=0$, and the residual deformation field at this level when $t=1$. We integrate $\phi^{(t)}$ from $t=0$ to $t=1$ by employing the scaling and squaring (ss) operation. Specifically, $\phi^{(1)}$ is achieved by repeating the loop: $\phi^{(\frac{1}{2^{t+1}})}=\phi^{(\frac{1}{2^{t}})}\circ\phi^{(\frac{1}{2^{t}})}$. After $T$ iterations, we obtain the final deformation field $\varphi$ at this level.
@@ 4.1 Datasets
To verify the efficacy of the proposed method, experiments were carried on three public datasets, including two brain MRI datasets and one abdomen CT dataset.
LPBA dataset (Shattuck et al., 2008) contains 40 brain MRI volumes, and each with 54 manually labeled region-of-interests (ROIs). All volumes have been rigidly pre-aligned to mni305. 30 volumes (30$\times$29 pairs) were employed for training and 10 volumes (10$\times$9 pairs) were used for testing.
Mindboggle dataset (Klein and Tourville, 2012) has been affinely aligned to mni152. Each brain MRI volume contains 62 manually labeled ROIs. 42 volumes (42$\times$41 pairs from the NKI-RS-22 and NKI-TRT-20 subsets) were employed for training, and 20 volumes from OASIS-TRT-20 (20$\times$19 pairs) were used for testing. All MRI volumes from the LPBA and Mindboggle datasets were pre-processed by min-max normalization, and skull-stripping using FreeSurfer (Fischl, 2012). The final size of each volume was 160$\times$192$\times$160 (1mm×$\times$1mm×$\times$1mm) after a center-cropping operation.
@@ 4.2 Evaluation Metrics
Accuracy evaluation. Dice score (DSC) (Dice, 1945) was computed as the primary similarity metric to evaluate the degree of overlap between corresponding regions. Moreover, the average symmetric surface distance (ASSD) (Taha and Hanbury, 2015) and the 95% maximum Hausdorff distance (HD95) (Huttenlocher et al., 1993) were calculated, which can reflect the similarity of the region contours. The quality of the predicted non-rigid deformation $\phi$ was assessed by the percentage of voxels with non-positive Jacobian determinant (i.e., folded voxels). All above metrics were calculated in 3D.
@@ 4.3 Comparison Methods
We compared our method with several cutting-edge registration methods: (1) SyN (Avants et al., 2008): a classical traditional approach, using the $SyNOnly$ setting in ANTS. (2) VoxelMorph(VM) (Balakrishnan et al., 2019): a popular CNN-based single-stage registration network. (3) VoxelMorph-diff(VM-diff) (Dalca et al., 2019): a popular CNN-based single-stage registration network. (4) TransMorph(TM) (Chen et al., 2022a): a single-stage registration network with SwinTransformer-enhanced encoder. (5) TransMatch (Chen et al., 2024): a Transformer-based registration network using cross attention modules for each level of the encoder and decoder. (6) PR++ (Kang et al., 2022): a pyramid registration network using 3D correlation layer. (7) DMR (Chen et al., 2022b): a pyramid registration network using Deformer module and multi-resolution refinement module. (8) PCnet (Lv et al., 2022): a pyramid network using several attention and weighting strategies to fuse multi-level feature maps and deformation fields. (9) PIViT (Ma et al., 2023): a light pyramid network using iterative SwinTransformer blocks in low resolution. (10) Im2grid(I2G) (Liu et al., 2022): a pyramid registration network with the coordinate Transformer. (11) ModeT (Wang et al., 2023): our conference version.
@@ 4.4 Implementation Details
We tested four different versions of our method, namely ModeTv2-s, ModeTv2-l, ModeTv2-diff-s, and ModeTv2-diff-l, which are the small model, large model, and their corresponding diffeomorphism versions with ss operation, respectively. The small model had $c$=8 convolutional channels in the first layer, as well as 6 channels for each head in the multi-head attention mechanism. The number of attention heads were set as 8, 4, 2, 1, 1 from bottom to top, respectively. This setting is the same as our conference version ModeT (Wang et al., 2023), allowing for easy comparison. Due to the implementation of CUDA operators, the large model version can be trained. Specifically, the large model had $c$=32 convolutional channels in the first layers, as well as 12 dimensions for each attention head. In the pyramid decoder, from coarse to fine, the number of attention heads were set as 32, 16, 8, 4, 1, respectively.
@@ Methods
HD95 (mm)
ASSD (mm)
%$\%|J_{\phi}|\leq 0$
Initial
45.7$\pm$10.3
33.53$\pm$9.89
12.26$\pm$4.15
SyN (Avants et al., 2008)
48.7$\pm$14.2
35.05$\pm$12.87
12.43$\pm$5.65
$<$0.3%
VM (Balakrishnan et al., 2019)
58.5$\pm$15.5
32.77$\pm$11.97
10.45$\pm$5.29
$<$3.%
VM-diff (Dalca et al., 2019)
51.2$\pm$14.3
35.32$\pm$13.43
12.13$\pm$5.63
$<$0.8%
TM (Chen et al., 2022a)
63.9$\pm$16.5
31.09$\pm$12.19
9.26$\pm$5.37
$<$2.%
TransMatch (Chen et al., 2024)
63.7$\pm$16.4
31.37$\pm$12.33
9.29$\pm$5.37
$<$3.%
PR++ (Kang et al., 2022)
68.8$\pm$15.9
29.22$\pm$12.98
8.08$\pm$5.26
$<$2.%
DMR (Chen et al., 2022b)
64.6$\pm$16.0
32.34$\pm$12.73
9.23$\pm$5.34
$<$3.%
PCnet (Lv et al., 2022)
72.4$\pm$15.5
27.00$\pm$12.91
7.10$\pm$5.10
$<$0.1%
PIViT (Ma et al., 2023)
71.2$\pm$14.1
26.92$\pm$12.19
7.07$\pm$4.08
$<$0.9%
I2G (Liu et al., 2022)
72.3$\pm$14.5
27.38$\pm$12.93
6.91$\pm$4.58
$<$0.5%
ModeT (Wang et al., 2023)
77.4$\pm$12.6
24.07$\pm$11.80
5.50$\pm$3.67
$<$0.9%
ModeTv2-s
78.5$\pm$12.7
23.61$\pm$12.47
5.44$\pm$3.83
$<$1.%
ModeTv2-diff-s
77.8$\pm$13.4
23.85$\pm$12.85
5.59$\pm$4.04
$<$0.01%
ModeTv2-l
80.6$\pm$11.7
21.92$\pm$12.12
4.95$\pm$3.49
$<$1.%
ModeTv2-diff-l 81.8$\pm$11.8 19.57$\pm$11.67 4.48$\pm$3.39 $<$0.1%
@@
Firstly, for VM-diff, $\sigma$ was set to 0.01, and $\lambda$ was set to 20 across all datasets. Then, regarding the remaining DL networks, the values of $\lambda$ were configured as follows: for the single-stage network on the LPBA dataset, $\lambda$ was set to 4, while for the pyramid network, it was set to 1. On both the Mindboggle and ABCT datasets, all networks were assigned $\lambda$ values of 1 and 0.5, respectively. For SyN, we used mean square error (MSE) as loss function. The neighborhood size $n$ was set as $3$. We used the Adam optimizer (Kingma and Ba, 2014) with a learning rate decay strategy as follows:
$lr_{m}=lr_{init}\cdot(1-\frac{m-1}{M})^{0.9},m=1,2,...,M$ (7)
where $lr_{m}$ represents the learning rate of $m$-th epoch and $lr_{init}$ represents the learning rate of initial epoch. For training the single-stage network on the LPBA dataset, we employed $lr_{init}=0.0004$. For all other training tasks, $lr_{init}=0.0001$ was applied. We set the batch size as $1$, $M$ as $30$ for the training of pyramid methods and $50$ for single-stage methods.
@@
HD95 (mm)
ASSD (mm)
%$\%|J_{\phi}|\leq 0$
Initial
53.7$\pm$4.8
7.43$\pm$0.78
2.74$\pm$0.33
SyN (Avants et al., 2008)
70.4$\pm$1.8
5.82$\pm$0.50
1.72$\pm$0.12
$<$0.002%
VM (Balakrishnan et al., 2019)
68.2$\pm$2.3
6.14$\pm$0.58
1.84$\pm$0.17
$<$0.004%
VM-diff (Dalca et al., 2019)
65.6$\pm$2.8
6.39$\pm$0.62
1.97$\pm$0.19
$<$0.0003%
TM (Chen et al., 2022a)
68.9$\pm$2.4
6.12$\pm$0.62
1.82$\pm$0.18
$<$0.02%
TransMatch (Chen et al., 2024)
67.8$\pm$2.5
6.20$\pm$0.61
1.87$\pm$0.18
$<$0.007%
PR++ (Kang et al., 2022)
69.5$\pm$2.2
6.11$\pm$0.59
1.76$\pm$0.17
$<$0.06%
DMR (Chen et al., 2022b)
69.2$\pm$2.4
6.14$\pm$0.62
1.79$\pm$0.18
$<$0.2%
PCnet (Lv et al., 2022)
72.0$\pm$1.9
5.71$\pm$0.55
1.62$\pm$0.14
$<$0.00006%
PIViT (Ma et al., 2023)
70.8$\pm$1.5
5.67$\pm$0.48
1.67$\pm$0.11
$<$0.02%
I2G (Liu et al., 2022)
71.0$\pm$1.4
5.69$\pm$0.47
1.64$\pm$0.10
$<$0.004%
ModeT (Wang et al., 2023)
72.1$\pm$1.4
5.54$\pm$0.47
1.58$\pm$0.11
$<$0.003%
ModeTv2-s
73.0$\pm$1.3
5.45$\pm$0.45
1.53$\pm$0.10
$<$0.001%
ModeTv2-diff-s
73.1$\pm$1.3
5.46$\pm$0.45
1.54$\pm$0.10
$<$0.000004%
ModeTv2-l 73.9$\pm$1.2 5.30$\pm$0.43 1.49$\pm$0.09 $<$0.008%
ModeTv2-diff-l
73.9$\pm$1.2
5.32$\pm$0.43
1.49$\pm$0.09
$<$0.000003%
@@
PO details on target domain. We conducted 50 iterations of fine-tuning on each model. For fine-tuning on brain data, $\lambda$ was set to 1, whereas abdominal data used $\lambda=0.5$. The learning rate was consistently maintained at $0.0001$ throughout the process. The PO experiments were performed using the Adam optimizer, while in a separate set of experiments, a comparison between the stochastic gradient descent (SGD) optimizer and the Adam optimizer was conducted.

### v2

@@ 3.3.1 GPU-accelerated ModeT
Let $F,M\in\mathbb{R}^{c\times h\times w\times l}$ be the fixed and moving feature maps from a specific layer of the hierarchical encoder, where $h,w,l$ are feature map dimensions, and $c$ is the channel count. These feature maps undergo linear projection ($proj$) and LayerNorm ($LN$) operations [4] to generate $Q$ (query) and $K$ (key):
@@
where the shared-weight projection operation originates from $N(0,1e^{-5})$, with biases initialized to 0. $Q$ and $K$ are then divided based on channels, resulting in $Q,K\in\mathbb{R}^{S\times h\times w\times l\times d}$, where $S$ represents the number of attention heads, and $d$ is the number of channels per attention head.
The neighborhood attention map is then calculated. Let $c(p)$ be the voxel $p$’s neighborhood, where $p\in\mathbb{R}^{3}$, representing the position of any voxel in $Q$. For a neighborhood of size $n\times n\times n$, $||c(p)||=n^{3}$. The multi-head neighborhood attention map is obtained through:
@@
where $B\in\mathbb{R}^{S\times n\times n\times n}$ is a learnable relative positional bias, initialized to zeros; and $NA(p,s)\in\mathbb{R}^{n\times n\times n}$. Zero padding is applied to the moving feature map to handle boundary voxels. Equation (3) shows how the neighborhood attention is computed for the $s$-th head at position $p$, breaking down semantic information from low resolution for individual similarity computation, laying the foundation for modeling diverse motion modes.
@@
where $\varphi^{(s)}\in\mathbb{R}^{h\times w\times l\times 3}$, $V\in\mathbb{R}^{n\times n\times n\times 3}$, and $V$ (value) represents the relative position coordinates for the neighborhood centroid. Specifically, $V_{x}=x-x_{c}$, where $x\in\mathbb{R}^{3}$ denotes a voxel’s position of $V$, and $x_{c}=\left(\lceil n/2\rceil,\lceil n/2\rceil,\lceil n/2\rceil\right)$ is the central voxel’s position of $V$. Importantly, $V$ remains unlearned, allowing the multi-head attention relationship to naturally transform into a multi-coordinate association.
@@ 3.3.2 Registration Head (RegHead)
We reevaluate the challenges associated with the diverse motion modes obtained through ModeT. In [69], a competitive weighting module (CWM) is designed to select a predominant motion pattern from various motion modes at a single point. However, weighting alone is insufficient to adequately address issues such as local field crossings that arise from multiple subfields (see Fig. 3(a)), since ModeT does not account for neighborhood relationships during calculation, or more precisely, only considers point to-point relationships. Simultaneously, to achieve a highly generalized network, we aim to minimize the number of trainable parameters on the deformation field estimator. We prefer the network to focus more on relearning the encoder’s feature extraction during PO.
@@
Under the condition of employing the diffeomorphic layer, we define the output of the convolution layer in RegHead as the stationary velocity field $v$ [16]. With regard to time $t$, the deformation field $\phi$ is defined as the ordinary differential equation (ODE):
@@
where $\phi^{(t)}$ is an identity transformation when $t=0$, and the residual deformation field at this level when $t=1$. We integrate $\phi^{(t)}$ from $t=0$ to $t=1$ by employing the scaling and squaring (ss) operation [1]. Specifically, $\phi^{(1)}$ is achieved by repeating the loop: $\phi^{(\frac{1}{2^{t+1}})}=\phi^{(\frac{1}{2^{t}})}\circ\phi^{(\frac{1}{2^{t}})}$. After $T$ iterations, we obtain the final deformation field $\varphi$ at this level.
It is worth noting that RegHead acts as a filter for motion patterns generated by ModeT. The convolutional layers within RegHead can learn kernels resembling mean or Gaussian filters to resolve most crossing artifacts. Furthermore, the ss layer refines the deformation field by performing multiple-step integration of velocity fields. This effectively decomposes large deformations into small diffeomorphic steps, preserving diffeomorphism over iterations.
@@ 4.1 Datasets
To verify the efficacy of the proposed method, experiments were carried on four public datasets, including three brain MRI datasets and one abdomen CT dataset.
LPBA dataset [57] contains 40 brain MRI volumes, and each with 54 manually labeled region-of-interests (ROIs). All volumes have been rigidly pre-aligned to mni305. 30 volumes (30$\times$29 pairs) were employed for training and 10 volumes (10$\times$9 pairs) were used for testing.
Mindboggle dataset [38] has been affinely aligned to mni152. Each brain MRI volume contains 62 manually labeled ROIs. 42 volumes (42$\times$41 pairs from the NKI-RS-22 and NKI-TRT-20 subsets) were employed for training, and 19 volumes from MMRR-21 (19$\times$18 pairs) were used for testing. All MRI volumes from the LPBA and Mindboggle datasets were pre-processed by min-max normalization, and skull-stripping using FreeSurfer [22]. The final size of each volume was 160$\times$192$\times$160 (1mm×$\times$1mm×$\times$1mm) after a center-cropping operation.
IXI dataset [^1] https://brain-development.org/ixi-dataset/, pre-processed by [9], includes brain MRI scans that have undergone affine pre-registration to the Talairach space, intensity normalization, and center cropping to dimensions of 160$\times$192$\times$224 (1mm×$\times$1mm×$\times$1mm). Each data contains 30 labeled ROIs. Following the protocol of [9], we conducted template-to-subject registration, utilizing 403 samples for training and 115 for testing. This task is typically employed for atlas-based segmentation.
@@ 4.2 Evaluation Metrics
Accuracy evaluation. Dice score (DSC) [17] was computed as the primary similarity metric to evaluate the degree of overlap between corresponding regions. Moreover, the average symmetric surface distance (ASSD) [63] and the 95% maximum Hausdorff distance (HD95) [33] were calculated, which can reflect the similarity of the region contours. The quality of the predicted non-rigid deformation $\phi$ was assessed by the percentage of voxels with non-positive Jacobian determinant (i.e., folded voxels). All above metrics were calculated in 3D.
@@ 4.3 Comparison Methods
We compared our method with several cutting-edge registration methods: (1) SyN [2]: a classical traditional approach, using the $SyNOnly$ setting in ANTS. (2) VoxelMorph(VM) [6]: a popular CNN-based single-stage registration network. (3) VoxelMorph-diff(VM-diff) [16]: a popular CNN-based single-stage registration network. (4) TransMorph(TM) [9]: a single-stage registration network with SwinTransformer-enhanced encoder. (5) TransMatch [14]: a Transformer-based registration network using cross attention modules for each level of the encoder and decoder. (6) PR++ [35]: a pyramid registration network using 3D correlation layer. (7) DeFormer [12]: a pyramid registration network using Deformer module and multi-resolution refinement module. (8) PCnet [44]: a pyramid network using several attention and weighting strategies to fuse multi-level feature maps and deformation fields. (9) PIViT [45]: a light pyramid network using iterative SwinTransformer blocks in low resolution. (10) Im2grid(I2G) [42]: a pyramid registration network with the coordinate Transformer. (11) ModeT [69]: our conference version.
@@ 4.4 Implementation Details
We tested four different versions of our method, namely ModeTv2-s, ModeTv2-l, ModeTv2-diff-s, and ModeTv2-diff-l, which are the small model, large model, and their corresponding diffeomorphism versions with ss operation, respectively. The small model had $c$=8 convolutional channels in the first layer, as well as 6 channels for each head in the multi-head attention mechanism. The number of attention heads were set as 8, 4, 2, 1, 1 from bottom to top, respectively. This setting is the same as our conference version ModeT [69], allowing for easy comparison. Due to the implementation of CUDA operators, the large model version can be trained. Specifically, the large model had $c$=32 convolutional channels in the first layers, as well as 12 dimensions for each attention head. In the pyramid decoder, from coarse to fine, the number of attention heads were set as 32, 16, 8, 4, 1, respectively.
Firstly, for VM-diff, $\sigma$ was set to 0.01, and $\lambda$ was set to 20 across all datasets. Then, regarding the remaining DL networks, the values of $\lambda$ were configured as follows: for the single-stage network on the LPBA dataset, $\lambda$ was set to 4, while for the pyramid network, it was set to 1. On the Mindboggle, ABCT, and IXI datasets, all networks were assigned $\lambda$ values of 1, 0.5, and 4, respectively. For SyN, we used mean square error (MSE) as loss function. The neighborhood size $n$ was set as $3$. We used the Adam optimizer [37] with a learning rate decay strategy as follows:
$lr_{m}=lr_{init}\cdot(1-\frac{m-1}{M})^{0.9},m=1,2,...,M$ (7)
where $lr_{m}$ represents the learning rate of $m$-th epoch and $lr_{init}$ represents the learning rate of initial epoch. For training the single-stage network on the LPBA dataset, we employed $lr_{init}=0.0004$. For the IXI dataset, $lr_{init}=0.0002$ was used. For all other training tasks, $lr_{init}=0.0001$ was applied. We set the batch size as $1$, $M$ as $30$ for the training of pyramid methods and $50$ for single-stage methods.
@@
Initial 45.7$\pm$10.3 33.53$\pm$9.89 12.26$\pm$4.15 -
SyN [2] 48.7$\pm$14.2 35.05$\pm$12.87 12.43$\pm$5.65 $<$0.2%
VM [6] 58.5$\pm$15.5 32.77$\pm$11.97 10.45$\pm$5.29 $<$11%
VM-diff [16] 51.2$\pm$14.3 35.32$\pm$13.43 12.13$\pm$5.63 $<$1%
TM [9] 63.9$\pm$16.5 31.09$\pm$12.19 9.26$\pm$5.37 $<$14%
TransMatch [14] 63.7$\pm$16.4 31.37$\pm$12.33 9.29$\pm$5.37 $<$13%
PR++ [35] 68.8$\pm$15.9 29.22$\pm$12.98 8.08$\pm$5.26 $<$6%
DeFormer [12] 64.6$\pm$16.0 32.34$\pm$12.73 9.23$\pm$5.34 $<$12%
PCnet [44] 72.4$\pm$15.5 27.00$\pm$12.91 7.10$\pm$5.10 $<$0.2%
PIViT [45] 71.2$\pm$14.1 26.92$\pm$12.19 7.07$\pm$4.08 $<$2%
I2G [42] 72.3$\pm$14.5 27.38$\pm$12.93 6.91$\pm$4.58 $<$0.9%
ModeT [69] 77.4$\pm$12.6 24.07$\pm$11.80 5.50$\pm$3.67 $<$3%
ModeTv2-s 78.5$\pm$12.7 23.61$\pm$12.47 5.44$\pm$3.83 $<$2%
ModeTv2-diff-s 77.8$\pm$13.4 23.85$\pm$12.85 5.59$\pm$4.04 $<$0.009%
ModeTv2-l 80.6$\pm$11.7 21.92$\pm$12.12 4.95$\pm$3.49 $<$4%
ModeTv2-diff-l 81.8$\pm$11.8 19.57$\pm$11.67 4.48$\pm$3.39 $<$0.2%
@@
Initial 53.7$\pm$4.8 7.43$\pm$0.78 2.74$\pm$0.33 -
SyN [2] 70.4$\pm$1.8 5.82$\pm$0.50 1.72$\pm$0.12 $<$0.0004%
VM [6] 68.2$\pm$2.3 6.14$\pm$0.58 1.84$\pm$0.17 $<$0.004%
VM-diff [16] 65.6$\pm$2.8 6.39$\pm$0.62 1.97$\pm$0.19 $<$0.0001%
TM [9] 68.9$\pm$2.4 6.12$\pm$0.62 1.82$\pm$0.18 $<$0.01%
TransMatch [14] 67.8$\pm$2.5 6.20$\pm$0.61 1.87$\pm$0.18 $<$0.009%
PR++ [35] 69.5$\pm$2.2 6.11$\pm$0.59 1.76$\pm$0.17 $<$0.2%
DeFormer [12] 69.2$\pm$2.4 6.14$\pm$0.62 1.79$\pm$0.18 $<$0.4%
PCnet [44] 72.0$\pm$1.9 5.71$\pm$0.55 1.62$\pm$0.14 $<$0.00003%
PIViT [45] 70.8$\pm$1.5 5.67$\pm$0.48 1.67$\pm$0.11 $<$0.04%
I2G [42] 71.0$\pm$1.4 5.69$\pm$0.47 1.64$\pm$0.10 $<$0.01%
ModeT [69] 72.1$\pm$1.4 5.54$\pm$0.47 1.58$\pm$0.11 $<$0.007%
ModeTv2-s 73.0$\pm$1.3 5.45$\pm$0.45 1.53$\pm$0.10 $<$0.002%
ModeTv2-diff-s 73.1$\pm$1.3 5.46$\pm$0.45 1.54$\pm$0.10 $<$0.0000007%
ModeTv2-l 73.9$\pm$1.2 5.30$\pm$0.43 1.49$\pm$0.09 $<$0.02%
ModeTv2-diff-l 73.9$\pm$1.2 5.32$\pm$0.43 1.49$\pm$0.09 $<$0.0000003%
@@
PO details on target domain. [^3] Note that iterative optimization was used only in the PO experiments (Section 4.5.3) to evaluate the cross-domain generalization capabilities of different network structures. For the accuracy experiments (Section 4.5.1) and convergence experiments (Section 4.5.2), no iterative optimization was employed, and the methodology aligns with typical DL approaches. We conducted 50 iterations of fine-tuning on each model. Note that considering IXI dataset was used for template-to-subject registration, while the other three datasets for inter-subject registration, we only employed LPBA, Mindboggle and ABCT for the PO experiments. For fine-tuning on brain data, $\lambda$ was set to 1, whereas abdominal data used $\lambda=0.5$. The learning rate was consistently maintained at $0.0001$ throughout the process. The PO experiments were performed using the Adam optimizer, while in a separate set of experiments, a comparison between the stochastic gradient descent (SGD) optimizer and the Adam optimizer was conducted.